1 Using the following to create a vertical analysis, what percentage would be reported for Property,...

1

Using the following to create a vertical analysis, what percentage would be reported for Property, Plant & Equipment?

2015 2014
Common Stock $ 300,000 $ 220,000
Current Assets $ 100,000 $ 110,000
Current Liabilities $    50,000 $    90,000
Intangible Assets $ 120,000 $ 130,000
Investments $ 200,000 $ 200,000
Long-term Liabilities $ 250,000 $ 275,000
Other Assets $    80,000 $    75,000
Property, Plant & Equipment $ 300,000 $ 250,000
Retained Earnings $ 200,000 $ 180,000
    32.7%
      62.3%
        25.6%
          44,2%

          2Use the following to determine total assets.

          Common Stock $300,000
          Current Assets $100,000
          Current Liabilities $50,000
          Intangible Assets $120,000
          Investments $200,000
          Long-term Liabilities $250,000
          Other Assets $80,000
          Property, Plant & Equipment $300,000
          Retained Earnings $200,000
          Answer #1

          1) Vertical analysis of property, plant and equipment :

          Total assets = 100000+120000+200000+80000+300000 = 800000

          2015 Vertical analysis = 300000/800000 = 37.5%

          2014 Vertical analysis = 250000/765000 = 32.7%

          So answer is a) 32.7%

          2) Total assets

          Current assets 100000
          Intangible assets 120000
          Investments 200000
          Other assets 80000
          Property, plant and equipment 300000
          Total assets 800000